# Single-photon scattering and bound states in a one-dimensional waveguide with topological giant atom

### v1

@@ Abstract
We investigate the single photon scattering and bound states in a coupled resonator waveguide (CRW) which couples to a topological giant atom (TGA) via two distant sites. Here, the TGA is constructed by a one dimensional Su-Schrieffer-Heeger chain with finite length. By modulating the topological phase of the TGA, the incident photon in the CRW can be completely reflected or transmitted, and is therefore beneficial to design the coherent photonic device. Meanwhile, we also achieve two pairs of bound states locating respectively above and blow the continuum. Whether the gap is open or closed depends on the boundary condition of the TGA. Therefore, the combination of the topology and the interference provides us an exciting opportunity to manipulate the photonic state in the context of waveguide QED.
@@ I INTRODUCTION
The light-matter interaction plays a crucial role in the fundamental sciences [1, 2], underpinning the rapid development of quantum technology. In recent years, the light-matter interaction in waveguide structures has attracted a lot of attentions, which have brought a lot of theoretical [3] and experimental studies in the waveguide QED [4] community, such as dressed or bound states [5, 6, 7, 8, 9, 10, 11], phase transitions [12, 13], single-photon devices [14, 15, 16], exotic topological and chiral phenomena [17, 18, 19, 20, 21, 22]. At the same time, the dynamical control of single-photon transmission is a hot topic in constructing quantum networks [23]. Since photons provide a reliable output of quantum information and the single-photon is considered to be one of the most suitable carriers of quantum information. In quantum devices and quantum networks [24, 25, 26], one-dimensional (1D) waveguides [27, 28] are essential light-matter interfaces, and the controllable single-photon transport with linear and nonlinear dispersion relationships has been extensively studied [11, 45, 54, 55, 56].
@@
In the traditional treatment, the light-matter interaction is often modeled by the dipole approximation, where the atoms are treated as point-like dipoles [29]. However, in recent years, the study of the interaction in atom-waveguide systems has been extended to the interaction between the photon and a giant atom, where the non-local light-matter interaction occurs with multiple points. In the giant atom community [30, 31, 32, 33, 34, 35, 36, 37], since the interactions between the waveguide photons and the giant atoms include multi-path quantum interferences, many new phenomena those do not exist in the traditional small atom systems have been predicted, such as tunable bound state [38, 39, 40, 41], decoherence free subspaces [42], electromagnetically induced transparency (EIT) [43, 44], Autler-Townes splitting (ATS) phenomena [45], chiral physics [46, 47, 48, 49, 50], and many others [51, 52, 53, 54, 55, 56]. The physical basis behind these phenomena is the interference and delay effects in the propagation of the photon/phonon between different coupling points.
@@
On the other hand, Haldane and Raghu [57, 58] proposed to manipulate the photon transport via topological structure in 2008, which paves the way for the development of topological photonics [59, 60, 61, 62]. In the simplest one dimensional case, the topological waveguide can be constructed [63, 64] through the Su-Schrieffer-Heeger (SSH) [65] chain, which is characterized by nonzero winding number or zero-mode edge state(s) in the topological nontrivial phase with periodical or open boundary condition, respectively.
Based on the above achievements, it is meaningful to combine the topological nature of the SSH chain and the interference effect from the giant atom, to manipulate the single photon state in the context of the waveguide QED. To this end, we construct a topological giant atom (TGA) by the SSH chain with finite length, which couples to the coupled resonator waveguide (CRW) via two distant sites. Our findings imply that the single photon scattering can be modulated by the topological phase of the TGA and the bound state is controlled by the boundary condition.
@@ II MODEL AND HAMILTONIAN
As schematically shown in Fig. 1, the system we consider is composed of two arrays of CRW. The upper one is a $N+1$ sites [i.e., $(N+1)/2$ pairs for odd $N$] SSH chain and the lower one is an infinite CRW with uniform hopping strength. Since the upper one supports a finite of discrete energy levels, we name it as the topological giant atom (TGA), whose topological nature is discussed below. The TGA couples to the CRW non-locally via two separate sites of $j=0$ and $j=N$. The Hamiltonian $H$ of the system can be divided into three parts, i.e., $H=H_{c}+H_{S}+H_{I}$.
Figure 1: Schematic diagram of coupling of a TGA to an infinite CRW via two coupling point.
@@
where $\omega_{c}$ is bare frequency of the resonators and $a_{j}^{\dagger}$ ($a_{j}$) is the bosonic creation (annihilation) operator on site $j$, $\xi$ is the hopping strength between the nearest resonators.
@@
In the TGA, each unit cell hosts $A$ and $B$ two resonators with the identical frequency $\omega_{e}$, and $C_{A,l}\left(C_{B,l}\right)$ is the annihilation operators at the $A(B)$ sublattices of the $l$th unit cell (see Fig. 1). The intra- and inter-cell hopping amplitudes are $t_{1}$ and $t_{2}$, respectively. Here, we use the value of $t_{3}$ to distinguish the boundary condition of the TGA. For the periodical boundary condition, we set $t_{3}=t_{2}$. In this case, the TGA is topologically nontrivial with a winding number of $w=1$ for $\left|t_{1}\right|<\left|t_{2}\right|$ and topologically trivial with $w=0$ for $\left|t_{1}\right|>\left|t_{2}\right|$ [66]. For the open boundary condition, we set $t_{3}=0$, and the TGA supports the boundary state in the topologically nontrivial phase, but not in the trivial phase.
The third part $H_{I}$ of the Hamiltonian describes the coupling between the TGA and the waveguide via the 00th and $N$th resonators with the same coupling strength $J$. Under the rotating wave approximation, the Hamiltonian $H_{I}$ can be written as
@@ III SINGLE-PHOTON SCATTERING
where $\left|G\right\rangle$ represents that all of the resonators in the waveguide are in the vacuum states, while the TGA is in the ground state $\left|g\right\rangle$. $U_{j}$ is the probability amplitude for finding a photonic excitation in the resonator $j$, and $X_{l}(Y_{l})$ describe the excitation amplitudes in the site $A_{l}(B_{l})$ of the TGA. In the regimes of $j<0$, $0\leqslant j\leqslant N$ and $j>N$, the amplitude $U_{j}$ can be written in the form
@@
where $r$ and $t$ are respectively the single-photon reflection and transmission amplitudes. Furthermore, the second line in the above equation implies that the incident photon can be transmitted (reflected) by the left(right) leg of the TGA giant atom with amplitude $A$ ($B$). The transmitted photon then propagates back and forth in the spatial regime covered by the giant atom with $0\leqslant j\leqslant N$.
@@
where $\Delta_{1}=E_{k}-\omega_{c}$ and $\Delta_{2}=E_{k}-\omega_{e}$ represent the detuning between the propagating photon in the waveguide and the bare resonators in the CRW and the TGA, respectively. Since the spontaneous radiation of the resonators and TGA are ignored, it satisfies that $R+T=1$, where $R=\left|r\right|^{2}$ is the reflection rate and $T=\left|t\right|^{2}$ is the transmission rate.
In Fig. 2 (c), we demonstrate the reflection rate $R$ as a function of the photon-atom detuning $\Delta_{2}$ in the resonance condition that is, $\omega_{c}=\omega_{e}$. The result for the scheme of the single cell [$N=1$ shown in Fig. 2 (a)] is illustrated by the blue solid curve. The two complete reflection peaks ($R=1$) characterize the inter coupling inside the TGA. Besides, when only one of the two sites in the TGA couple to the one resonator in the waveguide [shown in Fig. 2 (b)], the single photon reflection is characterized by the Rabi splitting behavior. As shown in the figure, the peaks of the orange dashed line locate exactly at the frequency $\Delta_{2}=\pm t_{1}$. The deviation from the Rabi splitting for the two sites coupling originates from the photonic interference effect when it propagates back and forth, and this interference is peculiar for the giant atom setup.
@@
Next, let us move to the case of $N>1$, in which the topological properties of the TGA take effect. In Fig. 3, we demonstrate the reflection rate $R$ for $N=5$ as a function of the detuning $\Delta_{2}$. As shown in the figures, the reflection rate shows a complicated behavior in the two bands regime of $-|t_{1}+t_{2}|<\Delta_{2}<-|t_{1}-t_{2}|$ and $|t_{1}-t_{2}|<\Delta_{2}<|t_{1}+t_{2}|$. In the gapped regime with $-|t_{1}-t_{2}|<\Delta_{2}<|t_{1}-t_{2}|$, the reflection rate behaves differently when the TGA undergoes the topological phase transition. For example, in the topological trivial phase ($t_{1}>t_{2}$), we illustrate the results for $t_{2}=0.1\xi$ in the cases of $t_{1}=0.2\xi$ and $t_{1}=0.5\xi$ respectively in Fig. 3(a). For small $t_{1}$ with $t_{1}=0.2\xi$, there exists a small peak at the zero detuning $\Delta_{2}=0$. As the increases of $t_{1}$, the peak is smoothed and it forms a relatively wide band for $R=0$ nearby the regime of $\Delta_{2}=0$. It implies that the TGA in the topological trivial phase can be used to realized the quantum cloaking [67, 68], that is, the TGA does not affect the propagation of the photon in the waveguide. On the other hand, when the TGA works in the topological non-trivial phase, the complete reflection window ($R=1$) is demonstrated in Fig. 3(b) nearby the resonant regime of $\Delta_{2}=0$. The results in Fig. 3 indicate that, we can use the TGA to construct the coherent single-photon device. When the value of $t_{1(2)}$ be is achieved by the order of $100$ MHz experimentally [69], we can realize the quantum cloaking or complete reflection with the bandwidth of about $100-200$ MHz, which is in the same order of $t_{1}$ and $t_{2}$.
@@ IV Single photon bound state
In the above section, we have studied the single photon scattering states who locate inside the propagating band $E_{k}\in\left[\omega_{c}-2\xi,\omega_{c}+2\xi\right]$. Due to the coupling between the TGA and the waveguide, the translational symmetry of the system is broken, leading to the photonic bound states which lay outside the propagating band. Here, we resort to the numerical diagonalization of the Hamiltonian to find these bound states when the TGA works in the topologically non-trivial phase ($t_{1}<t_{2}$).
@@
In Figs. 4 (a) and (b), we plot the energy spectrum by considering that the TGA works in the periodical boundary condition $(t_{3}=t_{2})$ and open boundary condition $(t_{3}=0)$, respectively. As expected, we observe the continual band which supports the propagating modes in both of the cases and the difference arises from the bound states outside the continuum. In the periodical boundary condition, there exist two pairs of non-degenerate bound states, which locate symmetrically above and below the continuum and are denoted by $E_{1},E_{2},E_{3}$ and $E_{4}$, respectively as shown in Fig. 4 (a). Within the available experimental parameter $\xi=100$ MHz, the gap between the gap state $\Delta_{E}=E_{1}-E_{2}=E_{3}-E_{4}$ is in the order of $10$MHz in the condition of $J=3\xi$. As a comparison, when the TGA works in the open boundary condition, the pairs of the bound states become degenerate as shown in Fig. 4 (b). Moreover, it satisfies $E_{5}=(E_{1}+E_{2})/2$ and $E_{6}=(E_{3}+E_{4})/2$.
@@
The above transition from the degenerate bound states to non-degenerate ones provide us an effective approach to detect the boundary condition of the TGA. To this end, we introduce an auxiliary probing atom whose resonant frequency satisfies $\omega_{p}=E_{6}$ and observe its excitation evolution. Then, the probing Hamiltonian is expressed as
@@
Preparing the probing atom in the excited state and the TGA-waveguide system in the ground state initially, we plot the excitation dynamics of the probing atom in Fig. 5 by taking $J=3\xi$. Here the probing atom is located in the resonator of $j=0$, that is, the left leg of the TGA. Since the probing atom is large detuned from both of the bound states in the condition of $f\ll|\omega-E_{3(4)}|$, the probing atom is effectively decoupled from the TGA-waveguide system in the open boundary condition. As a result, the dynamics of $P_{e}(t)=\langle|e\rangle_{p}\langle e|\rangle$ shows an exponential decay, which is determined by the spontaneous dissipation rate $\gamma$ as shown by the solid line in Fig. 5. In the periodical boundary condition, the pair of non-degenerate bound states emerge into the degenerate ones, which are resonate with the probing atom. So that, the dynamics exhibits a Rabi oscillation character as shown by the dashed line. In this sense, the boundary condition of the TGA can be detected in a coherent manner.

### v2

@@ Abstract
We investigate the single photon scattering and bound states in a coupled resonator waveguide (CRW) which couples to a topological giant atom (TGA) via two distant sites. Here, the TGA is constructed by a one dimensional Su-Schrieffer-Heeger (SSH) chain with finite length. By modulating the topological phase of the TGA, the incident photon in the CRW can be completely reflected or transmitted, and is therefore beneficial to design the coherent photonic device. Meanwhile, we also achieve two pairs of bound states locating respectively above and blow the continuum. Whether the gap is open or closed depends on the boundary condition of the TGA. Therefore, the combination of the topology and the interference provides us an exciting opportunity to manipulate the photonic state in the context of waveguide quantum electrodynamics.
@@ I INTRODUCTION
The light-matter interaction plays a crucial role in the fundamental sciences forp2019 ; gutr2021 , underpinning the rapid development of quantum technology. Recently, the light-matter interaction in waveguide structures has attracted a lot of attentions and brought a lot of theoretical zhay2021 and experimental studies in the waveguide quantum electrodynamics (QED) gux2017 community, such as dressed or bound states zheh2010 ; calg2016 ; shit2016 ; sane2017 ; fonp2017 ; calg2019 ; zhaw2020 , phase transitions fitm2017 ; qial2019 , single-photon devices shej2005 ; chad2007 ; zhol2008 , exotic topological and chiral phenomena rinm2014 ; yanv2014 ; miri2016 ; gong2016 ; belm2019 ; mahs2020 . At the same time, the dynamical control of single-photon transmission is a hot topic in constructing quantum networks kimh2008 . Since photons provide a reliable output of quantum information and the single-photon is considered to be one of the most suitable carriers of quantum information. In quantum devices and quantum networks lodp2017 ; royd2017 ; chad2018 , one-dimensional (1D) waveguides rits2012 ; lodp2018 are essential light-matter interfaces, and the controllable single-photon transport with linear and nonlinear dispersion relationships has been extensively studied caiq2021 ; zhangxj2023 ; lix2023 .
@@
In the traditional treatment, the light-matter interaction is often modeled by the dipole approximation, where the atoms are treated as point-like dipoles wald2008 . However, the study of the interaction in atom-waveguide systems has been extended to the interaction between the photon and giant atom(s) recently, where the non-local light-matter interaction occurs with multiple points. In the giant atom community pett2003 ; koca2014 ; koca2018 ; turp2019 ; andg2019 ; tuda2019 ; guos2020 ; Kanb2020 ; bagr2023 , since the interactions between the waveguide photons and the giant atoms include multi-path quantum interferences, many new phenomena those do not exist in the traditional small atom systems have been predicted, such as tunable bound state guol2020 ; lons2021 ; qiuy2023 ; limk2023 , decoherence free interaction cara2020 , electromagnetically induced transparency abda2022 ; vada2021 , Autler-Townes splitting phenomena zhaw2022 , chiral physics wanx2021 ; wanx2022 ; liun2022 ; Wagx2022 ; sora2022 , and many others dul2021 ; Dul2021 ; yuh2021 . The physical basis behind these phenomena is the interference and delay effects in the propagation of the photon/phonon between different coupling points.
@@
On the other hand, Haldane and Raghu half2008 ; rags2008 proposed to manipulate the photon transport via topological structure in 2008, which paved the way for the development of topological photonics blik2015 ; Blik2015 ; blik2019 ; ozat2019 ; jang1 ; jang2 ; jang3 ; jang4 . In the simplest one dimensional case, the topological waveguide can be constructed gux ; caiw2019 through the Su-Schrieffer-Heeger (SSH) suw1979 chain, which is characterized by nonzero winding number or zero-mode edge state(s) in the topological nontrivial phase with periodical boundary condition (PBC) or open boundary condition (OBC), respectively xuhs2022 ; jinl2017 ; jinl2019 ; wuhc2021 . Especially, the edge state in one dimensional SSH chain has been experimentally observed, and the localization length has been extracted in terms of the survival probability in Ref. wangl2022 .
Therefore, combining the unique effect of the nonlocal light matter interaction in giant atom quantum optics and the robust natural of the topological photonics, we here investigate the manipulation of single photon state in the context of waveguide QED. To this end, we couple a one dimensional SSH chain with finite length to an infinite CRW, to study the photonic scattering and dressing state in the single excitation level. Since the short topological SSH chain supports a few discrete energy levels and couples to the CRW via two distant site, we name it as topological giant atom (TGA) in what follows. Thus, the topological nature of the TGA is utilized to adjust the photon state in the CRW. Our findings imply that the single photon scattering can be modulated by the topological phase of the TGA and the bound state is controlled by its boundary conditions.
@@ II MODEL AND HAMILTONIAN
As schematically shown in Fig. 1, the system we consider is composed of two arrays of CRW. The upper one is a $N+1$ sites [i.e., $(N+1)/2$ pairs for odd $N$] SSH chain and the lower one is an infinite CRW with uniform photonic hopping strength. Since the upper SSH chain supports a finite of discrete energy levels and serves as TGA, whose topological nature is discussed below. The TGA couples to the CRW non-locally via two separate sites of $j=0$ and $j=N$. The Hamiltonian $H$ of the system can be divided into three parts, i.e., $H=H_{c}+H_{S}+H_{I}$.
Figure 1: Schematic diagram of coupling of a TGA to an infinite CRW via two coupling points. The upper part is a TGA formed by the finite SSH chain with $t_{1}$ and $t_{2}$ being intra- and inter-cell hopping amplitudes, respectively, and $t_{3}$ is the cyclic hopping amplitude. The lower part is an infinite CRW. The TGA coupled to the CRW via two separate sites of $j=0$ and $j=N$, and the coupling strengths for the two sites are both $J$.
@@
where $\omega_{c}$ is bare frequency of the resonators and $a_{j}^{\dagger}$ ($a_{j}$) is the bosonic creation (annihilation) operator on site $j$, and satisfies the commutation relation $[a_{j},a_{j}^{\dagger}]=1$. $\xi$ is the hopping strength between the nearest resonators.
@@
In the TGA, each unit cell hosts $A$ and $B$ two resonators with the identical frequency $\omega_{e}$, and $C_{A,l}\left(C_{B,l}\right)$ is the bosonic annihilation operator at the $A(B)$ sublattices of the $l$th unit cell (see Fig. 1). The intra- and inter-cell hopping amplitudes are $t_{1}$ and $t_{2}$, respectively. Here, we use the value of $t_{3}$ to distinguish the boundary condition of the TGA. For the PBC, we set $t_{3}=t_{2}$. In the case of PBC, the topological phase transition is characterized by the winding number $w$. In the topological trivial phase ($|t_{1}|>|t_{2}|$), the winding phase is $w=1$. Otherwise, in the the topological nontrivial phase ($|t_{1}|<|t_{2}|$), the winding phase is $w=0$. For the OBC, we set $t_{3}=0$, and the TGA supports the boundary state in the topologically nontrivial phase, but not in the trivial phase.
The third part $H_{I}$ of the Hamiltonian describes the coupling between the TGA and the CRW via the 00th and $N$th resonators with the same coupling strength $J$. Under the rotating wave approximation, the Hamiltonian $H_{I}$ can be written as
@@ III SINGLE-PHOTON SCATTERING
where $\left|G\right\rangle$ represents that all of the resonators in the CRW and TGA are in their vacuum states. $U_{j}$ is the probability amplitude for finding a photonic excitation in site $j$ of the CRW , and $X_{l}(Y_{l})$ describe the excitation amplitudes in site $A_{l}(B_{l})$ of the TGA. In the regimes of $j<0$, $0\leqslant j\leqslant N$ and $j>N$, the amplitude $U_{j}$ can be written in the form
@@
where $r$ and $t$ are respectively the single-photon reflection and transmission amplitudes. Furthermore, the second line in the above equation implies that the incident photon can be transmitted (reflected) by the left (right) leg of the TGA with amplitude $A$ ($B$). The transmitted photon then propagates back and forth in the spatial regime covered by the TGA with $0\leqslant j\leqslant N$.
@@
where $\Delta_{1}=E_{k}-\omega_{c}$ and $\Delta_{2}=E_{k}-\omega_{e}$ represent the detuning between the propagating photon in the waveguide and the bare resonators in the CRW and the TGA, respectively. Since the spontaneous radiation for all of the resonators are ignored, it satisfies that $R+T=1$, where $R=\left|r\right|^{2}$ is the reflection rate and $T=\left|t\right|^{2}$ is the transmission rate.
In Fig. 2 (c), we demonstrate the reflection rate $R$ as a function of the photon-atom detuning $\Delta_{2}$ in the resonance condition that is, $\omega_{c}=\omega_{e}$. The result for the scheme of the single cell [$N=1$ shown in Fig. 2 (a)] is illustrated by the blue solid curve. The two complete reflection peaks ($R=1$) characterize the inter coupling inside the TGA. Besides, when only one of the two sites in the TGA couple to one resonator in the lower CRW [shown in Fig. 2 (b)], the single photon reflection is characterized by the Rabi splitting behavior. As shown in the figure, the peaks of the orange dashed line locate exactly at the frequency $\Delta_{2}=\pm t_{1}$. The deviation from the Rabi splitting for the two sites coupling originates from the photonic interference effect when it propagates back and forth, and this interference is peculiar for the giant atom setup.
@@
Figure 4: The reflection rate $R$ as a function of detuning $\Delta_{2}$ when the bare resonator is off resonant with the TGA. The parameters are set as $J=0.9\xi$, $\omega_{e}=18\xi$, and $\omega_{c}=20\xi$.
Next, let us move to the case of $N>1$, in which the topological properties of the TGA take effect. In Fig. 3, we demonstrate the reflection rate $R$ for $N=5$ as a function of the detuning $\Delta_{2}$. As shown in the figures, the reflection rate shows a complicated behavior in the two bands regime of $-|t_{1}+t_{2}|<\Delta_{2}<-|t_{1}-t_{2}|$ and $|t_{1}-t_{2}|<\Delta_{2}<|t_{1}+t_{2}|$. In the gapped regime with $-|t_{1}-t_{2}|<\Delta_{2}<|t_{1}-t_{2}|$, the reflection rate behaves differently when the TGA undergoes the topological phase transition. For example, in the topological trivial phase ($t_{1}>t_{2}$), we illustrate the results for $t_{2}=0.1\xi$ in the cases of $t_{1}=0.2\xi$ and $t_{1}=0.5\xi$ respectively in Fig. 3(a). For small $t_{1}$ with $t_{1}=0.2\xi$, there exists a small peak at the zero detuning $\Delta_{2}=0$. As the increase of $t_{1}$, the peak is smoothed and it forms a relatively wide band for $R=0$ nearby the regime of $\Delta_{2}=0$. It implies that the incident photon will be completely transmitted. This is due to the constructive interference caused by the photonic back and forth propagation in the spatial regime covered by the TGA. It implies that the TGA in the topological trivial phase can be used to realize quantum cloaking fle2013 ; val2020 , that is, the presence of TGA has no effect on the propagation of the photon in the CRW. In other words, the photon in the CRW can not “see” the TGA, i.e., the TGA is invisible. On the other hand, when the TGA works in the topological non-trivial phase, the destructive interference between photons results in complete reflection and the reflection window ($R=1$) is demonstrated in Fig. 3(b) nearby the resonant regime of $\Delta_{2}=0$.
The results in Fig. 3 indicate that, we can use the TGA to construct the coherent single-photon device. When the value of $t_{1(2)}$ be is achieved by the order of $100$ MHz experimentally kim2021 , we can realize the quantum cloaking or complete reflection with the bandwidth of about $100-200$ MHz, which in the same order of $t_{1}$ and $t_{2}$.
We can continue to discuss the single photon scattering when the bare resonator is not in resonance with the TGA. In Fig. 4, we demonstrate reflectance $R$ as a function of photon-atom detuning $\Delta_{2}$ under non-resonant condition, i.e. $\omega_{e}\neq\omega_{c}$. The result of the single-cell scheme ($N=1$) is illustrated by the solid blue line. The dotted orange and yellow lines represent the result in the topologically trivial and non-trivial phase respectively for $N=5$. We find that the asymmetric reflection curve is completely different from the resonance condition, and around the reflection point, the reflection yields a Fano shape fano1961 ; fano2005 . With the increase of the size of the TGA, the reflection rate shows a complicated dependence on the detuning $\Delta_{2}$. This means that off resonance can induce Fano physics in such a hybrid system. We also observe that, there will be one or more completely reflection frequencies (excluding the edge of the photonic propagation band), depending on the values of $N$. It implies that, we can design on-demand single-photon transistors by adjusting the size of TGA in our waveguide QED setup.
@@ IV Single photon bound state
In the above section, we have studied the single photon scattering states which locate inside the propagating band $E_{k}\in\left[\omega_{c}-2\xi,\omega_{c}+2\xi\right]$. Due to the coupling between the TGA and the CRW, the translational symmetry of the system is broken, leading to the photonic bound states which lay outside the propagating band. Here, we resort to the numerical diagonalization of the Hamiltonian to find these bound states when the TGA works in the topologically non-trivial phase ($t_{1}<t_{2}$).
@@
In Figs. 5 (a) and (b), we plot the energy spectrum by considering that the TGA works in the PBC $(t_{3}=t_{2})$ and OBC $(t_{3}=0)$, respectively. As expected, we observe the continual band which supports the propagating modes in both of the cases and the difference arises from the bound states outside the continuum. In the PBC, there exist two pairs of non-degenerate bound states, which locate symmetrically above and below the continuum and are denoted by $E_{1},E_{2},E_{3}$ and $E_{4}$, respectively as shown in Fig. 5 (a). Within the available experimental parameter $\xi=100$ MHz, the gap between the gap state $\Delta_{E}=E_{1}-E_{2}=E_{3}-E_{4}$ is in the order of $10$ MHz in the condition of $J=3\xi$. As a comparison, when the TGA works in the OBC, the pairs of the bound states become degenerate as shown in Fig. 5 (b). Moreover, it satisfies $E_{5}=(E_{1}+E_{2})/2$ and $E_{6}=(E_{3}+E_{4})/2$.
@@
The above transition from the degenerate bound states to non-degenerate ones provides us an effective approach to detect the boundary condition of the TGA. To this end, we introduce an auxiliary probing atom whose resonant frequency satisfies $\omega_{p}=E_{6}$ and observe its excitation evolution. Then, the probing Hamiltonian is expressed as
@@
Preparing the probing atom in the excited state and the TGA-CRW coupled system in the ground state initially, we plot the excitation dynamics of the probing atom in Fig. 6 by taking $J=3\xi$. Here the probing atom is located in the resonator of $j=0$, that is, the left leg of the TGA. Since the probing atom is large detuned from both of the bound states in the condition of $f\ll|\omega-E_{3(4)}|$, the probing atom is effectively decoupled from the TGA-CRW system in the PBC. As a result, the dynamics of $P_{e}(t)=\langle|e\rangle_{p}\langle e|\rangle$ shows an exponential decay, which is determined by the spontaneous dissipation rate $\gamma$ as shown by the solid line in Fig. 6. In the OBC, the pair of non-degenerate bound states emerge into the degenerate ones, which resonate with the probing atom. So that, the dynamics exhibits a Rabi oscillation character as shown by the dashed line. In this sense, the boundary condition of the TGA can be detected in a coherent manner.